# A motivic Fundamental Lemma

### v3

@@ 5.4. Integration on generic fibers
In the case of the Hitchin fibration, the measure $\mu_{\mathrm{orb}}$ is given by a global volume form $\left\lvert\omega_{\mathrm{orb}}\right\rvert$, as explained in [22, Lemma 6.13]. For each $b\in B_{a}^{\flat}$, write $\omega_{b}$ for the volume forms on the smooth fibers $\widetilde{M}^{\square,t}_{G,b}$ and $\widetilde{M}^{\square,s}_{\widehat{G},b}$ obtained as a quotient of $\omega_{\mathrm{orb}}$ by a fixed volume form $\eta$ on the Hitchin base $\widetilde{{\mathcal{A}}}^{{\mathrm{ani}}}$. Proposition 2.9.4, a version of Fubini’s theorem, shows that
@@
From its definition, $\chi_{s}$ takes a finite number of values, say indexed by a finite group $\mu$. Since it is a constructible function, there is a $b$-definable partition of $\widetilde{{\mathbf{M}}}^{\square}_{G,b}$ into parts $(X_{g})_{g\in\mu}$ such that on $X_{g}$, the function $\chi_{s}$ is equal to $g$. The statement of the lemma is then equivalent to the truth of a first-order sentence $\phi(b)$ with parameter $b$. We then need to show that $\phi(b)$ holds for generic $b$. By Propositions 2.10.1 and 4.8.4, the specialization $\phi_{L}(b)$ of $\phi(b)$ to a local field $L$ of large enough residue characteristic is equivalent to the equivalent lemma for an instance of the Geometric Stabilization theorem over a reductive group over $L$. For generic $b$, the fact that $\phi_{L}(b)$ holds is the key ingredient in the proof of Lemma 6.14 in [22]. By Proposition 2.10.1, we conclude that $\phi(b)$ holds for generic $b$. ∎
@@
Similarly to Lemma 5.5.2, the multiplicativity of $\chi_{s}\circ e$ is implied by the multiplicativity of its specialization to local fields, where it follows from a reinterpretation $\chi_{s}\circ e$ in terms of Tate duality, see [22, Section 6.5]. Since the Néron model is not in general definable in families, we cannot readily deduce the factorisation property. However, using multiplicativity, it suffices to show that $\chi_{s}\circ e$ is constant to 1 on the neutral component of the Néron model. Assume not, let $x$ in the neutral component of the Néron model be such that $\chi_{s}\circ e(x)\neq 1$. Since elements in the neutral component of a Néron model are divisible, up to replacing $K^{\prime}$ by a finite extension $K^{\prime}_{N}$, we can assume that there exists $y\in\widetilde{{\mathcal{P}}}^{\square}_{G,b}(K^{\prime}_{N}\mathopen{(\!(}t\mathopen{)\!)})$ such that $y^{N}=x$, where $N$ is the order of the image of $\chi_{s}\circ e$, and the extension $K^{\prime}_{N}$ is of degree depending only on $N$. Since this extension is pseudo-finite, the function $\chi_{s}\circ e$ is defined as well on this extension, multiplicative, and compatible with the restriction to $K^{\prime}$. By compatibility of the Néron model with unramified base change, $x$ lies in the neutral component of the Néron model of $\widetilde{{\mathcal{P}}}^{\square}_{G,b}(K^{\prime}_{N}\mathopen{(\!(}t\mathopen{)\!)})$. Hence we get a contradiction, since if $\chi_{s}\circ e(x)\neq 1$, then $\chi_{s}\circ e(y)$ is of order more than $N$. ∎
@@
$\int^{\mathrm{mot}}_{\widetilde{{\mathbf{M}}}^{\square,t}_{G,b}}\chi_{s}\circ e\left\lvert\omega_{b}\right\rvert=\begin{cases}0\ \ \ \text{ if }i(t)\neq 0\text{ or }i(s)\neq 0\\ \chi_{\mathrm{Psf},\mathrm{rel}}(\overline{{\mathcal{N}}}(\widetilde{{\mathbf{P}}}^{\square}_{G,b})\mathbb{L}^{-\operatorname{ord}(\omega_{b})})\ \ \ \text{if }i(t)=0\text{ and }i(s)=0.\end{cases}$
@@
The last equation follows since the volume of an abelian variety with respect to any volume form equals the class of the special fiber of the Néron model up to a shift by the order of vanishing of the volume form along the special fiber, see for example [33] or [34, Remark 4.1.1].
@@
We can now finish the proof of Theorem 5.4.1. By Proposition 5.5.4, if either $i(t)\neq 0$ or $i(s)\neq 0$, then both sides are zero hence the equality holds. In the case $i(t)=i(s)=0$, we need to prove that $\chi_{\mathrm{Psf},\mathrm{rel}}(\overline{{\mathcal{N}}}(\widetilde{{\mathbf{P}}}^{\square}_{G,b}))=\chi_{\mathrm{Psf},\mathrm{rel}}(\overline{{\mathcal{N}}}(\widetilde{{\mathbf{P}}}^{\square}_{\widehat{G},b}))$. By the main result of [15], $\widetilde{{\mathcal{P}}}^{\square}_{\widehat{G},b}$ and ${\widetilde{{\mathcal{P}}}^{\square}}_{G,b}$ are dual abelian varieties. Since the Grothendieck pairing on Néron models is non-degenerate in residue-characteristic 00 [23, 11.3, Exposé IX] and $K^{\prime}$ is pseudo-finite we deduce
@@
as in the proof of [35, Proposition 4.3]. Finally, the isogeny between $\widetilde{{\mathcal{P}}}^{\square}_{G,b}$ and $\widetilde{{\mathcal{P}}}^{\square}_{\widehat{G},b}$ induces an isogeny between their Néron models by [5, Proposition 7.3.6]. The motives of two isogenous connected commutative groups are equal by [1, Theorem 3.3 (4)] and [34, Lemma 2.5.1]. The motive $\chi_{\mathrm{Psf},\mathrm{rel}}(\overline{{\mathcal{N}}}(\widetilde{{\mathbf{P}}}^{\square}_{G,b}))$ is $|\phi_{{\widetilde{{\mathcal{P}}}^{\square}}_{G,b}}|$ times the motive of the neutral component, hence the motives $\chi_{\mathrm{Psf},\mathrm{rel}}(\overline{{\mathcal{N}}}(\widetilde{{\mathbf{P}}}^{\square}_{G,b}))$ and $\chi_{\mathrm{Psf},\mathrm{rel}}(\overline{{\mathcal{N}}}(\widetilde{{\mathbf{P}}}^{\square}_{\widehat{G},b}))$ are equal, which concludes the proof of Theorem 5.4.1.
@@ 6.1. Statement
Let $H$ be an endoscopic group of $G$, ${\mathbf{H}}$ the associated definable group and ${\mathbf{h}}$ the Lie algebra. We add subscripts $H$ to the various objects associated to $H$, such as ${\mathbf{c}}_{H}$. We have an inclusion ${\mathbf{c}}_{H}\to{\mathbf{c}}$. Let $a_{H}\in{\mathbf{c}}_{H}(L)$ be regular semi-simple, and $a$ its image in ${\mathbf{c}}(L)$. The group $H$ determines a character $\kappa$ of $\mathrm{H}^{1}(L,J_{a})$, where $J_{a}\cong I_{\varepsilon(a)}$ denotes the regular centralizer of $a$, see [37, Section 1.4]. Notice that $J_{a}(L)=T(L)$ by definition.
@@
Recall from Ngô [37, Théorème 1] the statement of the Fundamental Lemma of Langlands-Shelstad for Lie algebras.
@@ 6.2. From orbital integrals to affine Springer fibers
Here the right hand side can be identified with the $K(a)$-points of ${\mathbf{\Lambda}}\backslash{\mathcal{N}}_{a}$, where ${\mathcal{N}}_{a}$ denotes the affine Springer fiber associated with $a$. Since it is not of finite type, ${\mathcal{N}}_{a}$ is not definable. However, by [31], or more precisely [37, Proposition 3.4.1], the quotient ${\mathbf{\Lambda}}\backslash{\mathcal{N}}_{a}$ is a projective finite type scheme whose points can be encoded by a definable set. We thus get
@@
Now let $\gamma^{\prime}\in{\mathbf{g}}$ be stably conjugate to $\varepsilon(a)$ and let $\delta=\operatorname{\mathsf{inv}}(\varepsilon(a),\gamma^{\prime})\in\mathrm{H}^{1}(L,J_{a})$ denote the corresponding class. By [37, Lemme 8.2.4] we have an isomorphism
@@
where $J^{\prime}_{a}\to J_{a}$ is a group scheme with connected fibers generically isomorphic to $J_{a}$ and ${\mathbb{P}}(J^{\prime}_{a})$ denotes the Picard-stack of $J^{\prime}_{a}$-torsors with a generic trivialization. Since $J^{\prime}_{a}$ is connected, the $K$-points of ${\mathbb{P}}(J^{\prime}_{a})$ can be identified with $J_{a}(L)/J^{\prime}_{a}({\mathcal{O}}_{L})$, see [37, Section 3.3]. The lattice ${\mathbf{\Lambda}}$ thus naturally embeds into ${\mathbb{P}}(J^{\prime}_{a})$ and the quotient ${\mathbf{\Lambda}}\backslash{\mathbb{P}}(J^{\prime}_{a})$ is an affine finite type group scheme [37, Lemme 3.8.1]. We write ${\mathbf{\Lambda}}\backslash{\mathbb{P}}({\mathbf{J}}^{\prime}_{a})$ for the associated definable set and $\pi_{0}(a)$ for its group of connected components. Since ${\mathbf{\Lambda}}$ is discrete and torsion free we have by Lang’s theorem an inclusion
@@
where the $\kappa$-isotypical component is defined as in Section 4.4 and $\mathcal{D}$ is identified with a subgroup of $\mathrm{H}^{1}(L,J_{a})\cong\mathrm{H}^{1}(K,{\mathbb{P}}(J^{0}_{a}))$ through $\operatorname{\mathsf{inv}}(\varepsilon(a),\cdot)$ as in Section 6.1. Here we also used [37, Lemme 8.2.6], which implies that $({\mathbf{\Lambda}}\backslash{\mathcal{N}}_{a})^{\delta}$ has no $K$-rational point for $\delta\notin\mathcal{D}$ and thus $\chi_{\mathrm{Psf}}(({\mathbf{\Lambda}}\backslash{\mathcal{N}}_{a})^{\delta})=0$ in this case.
@@
[…]
@@
which is the motivic analogue of [37, Proposition 8.2.5]. Notice the equality $\operatorname{\mathsf{vol}}({\mathbf{J}}^{\prime}_{a}({\mathcal{O}}),\omega_{T})=\mathbb{L}^{\alpha}\chi_{\mathrm{Psf}}(({\mathbf{J}}^{\prime}_{a})_{K})$, with $\alpha$ depending on the choice of $\omega_{T}$, in particular $\operatorname{\mathsf{vol}}({\mathbf{J}}^{\prime}_{a}({\mathcal{O}}),\omega_{T})$ is the motive of a commutative algebraic group.
@@ 6.3. From affine Springer to Hitchin fibers
The connection between Springer and Hitchin fibers is given by the product formula [37, Proposition 4.15.1]. Consider the geometric setup of $G$-Higgs bundles on a curve $X$ in the generality of Section 4.7. Let $a\in\operatorname{\mathsf{A}}^{{\mathrm{ani}}}$ be a point with pseudo-finite residue field. By definition we get a morphism $a:X\to\mathfrak{c}_{D}$ and we denote by $U\subseteq X$ the open dense preimage of $\mathfrak{c}_{D}^{rs}$ under $a$. For any closed point $\nu\in X\setminus U$ the completion of the local ring at $\nu$ is a discretely valued field $L_{\nu}$ with pseudo-finite residue field $K_{\nu}$. We write a subscript $\nu$ for the restriction of $G$, $T$, $a$ etc. to this formal neighbourhood after having fixed a trivialization of $D$. In particular $a_{\nu}\in\mathfrak{c}_{\nu}^{rs}(L_{\nu})$ gives rise to an affine Springer fiber ${\mathcal{N}}_{a_{\nu}}$.
@@
and similarly for the symmetry groups $\prod_{\nu\in X\setminus U}{\mathbb{P}}(J_{a_{\nu}})\rightarrow{\mathbb{P}}_{a}$. The product formula [37, Proposition 4.15.1], see also [37, Proposition 8.4.2], states that over $\overline{K}$ we have a homeomorphism of proper DM-stacks
@@
compatible with the actions of $\operatorname{Gal}(\overline{K}/K)$ and ${\mathbb{P}}_{a}$. Furthermore let ${\mathbb{P}}^{\prime}_{a}\to{\mathbb{P}}(J_{a})$ be as in [37, Section 8.4], in particular ${\mathbb{P}}^{\prime}_{a}$ is an algebraic group locally of finite type. It follows essentially by definition that (6.3.1) continues to hold with ${\mathbb{P}}(J_{a_{\nu}})$ and ${\mathbb{P}}_{a}$ replaced by ${\mathbb{P}}(J^{\prime}_{a_{\nu}})$ and ${\mathbb{P}}^{\prime}_{a}$, or more conveniently for us, we have a fibration
@@
The formula is the motivic analogue of [37, Corollaire 8.4.4].

### v4

@@ 5.4. Integration on generic fibers
In the case of the Hitchin fibration, the measure $\mu_{\mathrm{orb}}$ is given by a global volume form $\left\lvert\omega_{\mathrm{orb}}\right\rvert$, as explained in [23, Lemma 6.13]. For each $b\in B_{a}^{\flat}$, write $\omega_{b}$ for the volume forms on the smooth fibers $\widetilde{M}^{\square,t}_{G,b}$ and $\widetilde{M}^{\square,s}_{\widehat{G},b}$ obtained as a quotient of $\omega_{\mathrm{orb}}$ by a fixed volume form $\eta$ on the Hitchin base $\widetilde{{\mathcal{A}}}^{{\mathrm{ani}}}$. Proposition 2.9.4, a version of Fubini’s theorem, shows that
@@
From its definition, $\chi_{s}$ takes a finite number of values, say indexed by a finite group $\mu$. Since it is a constructible function, there is a $b$-definable partition of $\widetilde{{\mathbf{M}}}^{\square}_{G,b}$ into parts $(X_{g})_{g\in\mu}$ such that on $X_{g}$, the function $\chi_{s}$ is equal to $g$. The statement of the lemma is then equivalent to the truth of a first-order sentence $\phi(b)$ with parameter $b$. We then need to show that $\phi(b)$ holds for generic $b$. By Propositions 2.10.1 and 4.9.4, the specialization $\phi_{L}(b)$ of $\phi(b)$ to a local field $L$ of large enough residue characteristic is equivalent to the equivalent lemma for an instance of the Geometric Stabilization theorem over a reductive group over $L$. For generic $b$, the fact that $\phi_{L}(b)$ holds is the key ingredient in the proof of Lemma 6.14 in [23]. By Proposition 2.10.1, we conclude that $\phi(b)$ holds for generic $b$. ∎
@@
Similarly to Lemma 5.5.2, the multiplicativity of $\chi_{s}\circ e$ is implied by the multiplicativity of its specialization to local fields, where it follows from a reinterpretation $\chi_{s}\circ e$ in terms of Tate duality, see [23, Section 6.5]. Since the Néron model is not in general definable in families, we cannot readily deduce the factorisation property. However, using multiplicativity, it suffices to show that $\chi_{s}\circ e$ is constant to 1 on the neutral component of the Néron model. Assume not, let $x$ in the neutral component of the Néron model be such that $\chi_{s}\circ e(x)\neq 1$. Since elements in the neutral component of a Néron model are divisible, up to replacing $K^{\prime}$ by a finite extension $K^{\prime}_{N}$, we can assume that there exists $y\in\widetilde{{\mathcal{P}}}^{\square}_{G,b}(K^{\prime}_{N}\mathopen{(\!(}t\mathopen{)\!)})$ such that $y^{N}=x$, where $N$ is the order of the image of $\chi_{s}\circ e$, and the extension $K^{\prime}_{N}$ is of degree depending only on $N$. Since this extension is pseudo-finite, the function $\chi_{s}\circ e$ is defined as well on this extension, multiplicative, and compatible with the restriction to $K^{\prime}$. By compatibility of the Néron model with unramified base change, $x$ lies in the neutral component of the Néron model of $\widetilde{{\mathcal{P}}}^{\square}_{G,b}(K^{\prime}_{N}\mathopen{(\!(}t\mathopen{)\!)})$. Hence we get a contradiction, since if $\chi_{s}\circ e(x)\neq 1$, then $\chi_{s}\circ e(y)$ is of order more than $N$. ∎
@@
[…]
@@
The last equation follows since the volume of an abelian variety with respect to any volume form equals the class of the special fiber of the Néron model up to a shift by the order of vanishing of the volume form along the special fiber, see for example [34] or [35, Remark 4.1.1].
@@
We can now finish the proof of Theorem 5.4.1. By Proposition 5.5.4, if either $i(t)\neq 0$ or $i(s)\neq 0$, then both sides are zero hence the equality holds. In the case $i(t)=i(s)=0$, we need to prove that $\chi_{\mathrm{Psf},\mathrm{rel}}(\overline{{\mathcal{N}}}(\widetilde{{\mathbf{P}}}^{\square}_{G,b}))=\chi_{\mathrm{Psf},\mathrm{rel}}(\overline{{\mathcal{N}}}(\widetilde{{\mathbf{P}}}^{\square}_{\widehat{G},b}))$. By the main result of [15], $\widetilde{{\mathcal{P}}}^{\square}_{\widehat{G},b}$ and ${\widetilde{{\mathcal{P}}}^{\square}}_{G,b}$ are dual abelian varieties. Since the Grothendieck pairing on Néron models is non-degenerate in residue-characteristic 0 [24, 11.3, Exposé IX] and $K^{\prime}$ is pseudo-finite we deduce
@@
as in the proof of [36, Proposition 4.3]. Finally, the isogeny between $\widetilde{{\mathcal{P}}}^{\square}_{G,b}$ and $\widetilde{{\mathcal{P}}}^{\square}_{\widehat{G},b}$ induces an isogeny between their Néron models by [5, Proposition 7.3.6]. The motives of two isogenous connected commutative groups are equal by [1, Theorem 3.3 (4)] and [35, Lemma 2.5.1]. The motive $\chi_{\mathrm{Psf},\mathrm{rel}}(\overline{{\mathcal{N}}}(\widetilde{{\mathbf{P}}}^{\square}_{G,b}))$ is $|\phi_{{\widetilde{{\mathcal{P}}}^{\square}}_{G,b}}|$ times the motive of the neutral component, hence the motives $\chi_{\mathrm{Psf},\mathrm{rel}}(\overline{{\mathcal{N}}}(\widetilde{{\mathbf{P}}}^{\square}_{G,b}))$ and $\chi_{\mathrm{Psf},\mathrm{rel}}(\overline{{\mathcal{N}}}(\widetilde{{\mathbf{P}}}^{\square}_{\widehat{G},b}))$ are equal, which concludes the proof of Theorem 5.4.1.
@@ 6.1. Statement
Let $H$ be an endoscopic group of $G$, $\mathbf{H}$} the associated definable group and $\mathbf{h}$} the Lie algebra. We add subscripts $H$ to the various objects associated to $H$, such as ${\mathbf{c}}_{H}$. We have an inclusion ${\mathbf{c}}_{H}\to{\mathbf{c}}$. Let $a_{H}\in{\mathbf{c}}_{H}(L)$ be regular semi-simple, and $a$ its image in ${\mathbf{c}}(L)$. The group $H$ determines a character $\kappa$ of $\mathrm{H}^{1}(L,J_{a})$, where $J_{a}\cong I_{\varepsilon(a)}$ denotes the regular centralizer of $a$, see [38, Section 1.4]. Notice that $J_{a}(L)=T(L)$ by definition.
@@
Recall from Ngô [38, Théorème 1] the statement of the Fundamental Lemma of Langlands-Shelstad for Lie algebras.
@@ 6.2. From orbital integrals to affine Springer fibers
Here the right hand side can be identified with the $K(a)$-points of ${\mathbf{\Lambda}}\backslash{\mathcal{N}}_{a}$, where ${\mathcal{N}}_{a}$ denotes the affine Springer fiber associated with $a$. Since it is not of finite type, ${\mathcal{N}}_{a}$ is not definable. However, by [32], or more precisely [38, Proposition 3.4.1], the quotient ${\mathbf{\Lambda}}\backslash{\mathcal{N}}_{a}$ is a projective finite type scheme whose points can be encoded by a definable set. We thus get
@@
Now let $\gamma^{\prime}\in{\mathbf{g}}$ be stably conjugate to $\varepsilon(a)$ and let $\delta=\operatorname{\mathsf{inv}}(\varepsilon(a),\gamma^{\prime})\in\mathrm{H}^{1}(L,J_{a})$ denote the corresponding class. By [38, Lemme 8.2.4] we have an isomorphism
@@
where $J^{\prime}_{a}\to J_{a}$ is a group scheme with connected fibers generically isomorphic to $J_{a}$ and ${\mathbb{P}}(J^{\prime}_{a})$ denotes the Picard-stack of $J^{\prime}_{a}$-torsors with a generic trivialization. Since $J^{\prime}_{a}$ is connected, the $K$-points of ${\mathbb{P}}(J^{\prime}_{a})$ can be identified with $J_{a}(L)/J^{\prime}_{a}({\mathcal{O}}_{L})$, see [38, Section 3.3]. The lattice $\mathbf{\Lambda}$} thus naturally embeds into ${\mathbb{P}}(J^{\prime}_{a})$ and the quotient ${\mathbf{\Lambda}}\backslash{\mathbb{P}}(J^{\prime}_{a})$ is an affine finite type group scheme [38, Lemme 3.8.1]. We write ${\mathbf{\Lambda}}\backslash{\mathbb{P}}({\mathbf{J}}^{\prime}_{a})$ for the associated definable set and $\pi_{0}(a)$ for its group of connected components. Since $\mathbf{\Lambda}$} is discrete and torsion free we have by Lang’s theorem an inclusion
@@
where the $\kappa$-isotypical component is defined as in Section 4.4 and $\mathcal{D}$ is identified with a subgroup of $\mathrm{H}^{1}(L,J_{a})\cong\mathrm{H}^{1}(K,{\mathbb{P}}(J^{0}_{a}))$ through $\operatorname{\mathsf{inv}}(\varepsilon(a),\cdot)$ as in Section 6.1. Here we also used [38, Lemme 8.2.6], which implies that $({\mathbf{\Lambda}}\backslash{\mathcal{N}}_{a})^{\delta}$ has no $K$-rational point for $\delta\notin\mathcal{D}$ and thus $\chi_{\mathrm{Psf}}(({\mathbf{\Lambda}}\backslash{\mathcal{N}}_{a})^{\delta})=0$ in this case.
@@
$0\longrightarrow\pi_{0}(a)(K)\longrightarrow\pi_{0}(a)(K^{\mathrm{alg}})\xrightarrow{\>1-\sigma\>}\pi_{0}(a)(K^{\mathrm{alg}})\longrightarrow\mathrm{H}^{1}(K,\pi_{0}(a))\longrightarrow 0,$
@@
which is the motivic analogue of [38, Proposition 8.2.5]. Notice the equality $\operatorname{\mathsf{vol}}({\mathbf{J}}^{\prime}_{a}({\mathcal{O}}),\omega_{T})=\mathbb{L}^{\alpha}\chi_{\mathrm{Psf}}(({\mathbf{J}}^{\prime}_{a})_{K})$, with $\alpha$ depending on the choice of $\omega_{T}$, in particular $\operatorname{\mathsf{vol}}({\mathbf{J}}^{\prime}_{a}({\mathcal{O}}),\omega_{T})$ is the motive of a commutative algebraic group.
@@ 6.3. From affine Springer to Hitchin fibers
The connection between Springer and Hitchin fibers is given by the product formula [38, Proposition 4.15.1]. Consider the geometric setup of $G$-Higgs bundles on a curve $X$ in the generality of Section 4.7. Let $a\in\operatorname{\mathsf{A}}^{{\mathrm{ani}}}$ be a point with pseudo-finite residue field. By definition we get a morphism $a:X\to\mathfrak{c}_{D}$ and we denote by $U\subseteq X$ the open dense preimage of $\mathfrak{c}_{D}^{rs}$ under $a$. For any closed point $\nu\in X\setminus U$ the completion of the local ring at $\nu$ is a discretely valued field $L_{\nu}$ with pseudo-finite residue field $K_{\nu}$. We write a subscript $\nu$ for the restriction of $G$, $T$, $a$ etc. to this formal neighbourhood after having fixed a trivialization of $D$. In particular $a_{\nu}\in\mathfrak{c}_{\nu}^{rs}(L_{\nu})$ gives rise to an affine Springer fiber ${\mathcal{N}}_{a_{\nu}}$.
@@
and similarly for the symmetry groups $\prod_{\nu\in X\setminus U}{\mathbb{P}}(J_{a_{\nu}})\rightarrow{\mathbb{P}}_{a}$. The product formula [38, Proposition 4.15.1], see also [38, Proposition 8.4.2], states that over $\overline{K}$ we have a homeomorphism of proper DM-stacks
@@
compatible with the actions of $\operatorname{Gal}(\overline{K}/K)$ and ${\mathbb{P}}_{a}$. Furthermore let ${\mathbb{P}}^{\prime}_{a}\to{\mathbb{P}}(J_{a})$ be as in [38, Section 8.4], in particular ${\mathbb{P}}^{\prime}_{a}$ is an algebraic group locally of finite type. It follows essentially by definition that (6.3.1) continues to hold with ${\mathbb{P}}(J_{a_{\nu}})$ and ${\mathbb{P}}_{a}$ replaced by ${\mathbb{P}}(J^{\prime}_{a_{\nu}})$ and ${\mathbb{P}}^{\prime}_{a}$, or more conveniently for us, we have a fibration
@@
The formula is the motivic analogue of [38, Corollaire 8.4.4].